[figure: 9.7 cm]

MAGNETIC PROPERTIES OF BULK AMORPHOUS Nd_{50}Fe_{40}Si_{10−x}Al_{x} ALLOYS

H. Chiriac and N. Lupu National Institute of R&D for Technical Physics, 6600 Iasi, Romania

Some bulk amorphous alloys of compositions Nd_{50}Fe_{40}Si_{10−x}Al_{x} (x between 0 and 10 at. %) were investigated with the view to enlarge the supercooled liquid region and to improve their hard magnetic properties. The increase of the Al content leads to the increase of the glass-forming ability and allows to obtain amorphous bars up to 3 mm in diameter by mould casting method. Coercivities up to 315 kA/m and saturation magnetization up to 0.33 T were measured for the Nd_{50}Fe_{40}Si_{10} thick amorphous ribbons prepared by melt-spinning. For amorphous ribbons, independent of their thickness, the substitution of Si by Al does not change the saturation magnetization, but causes a decrease of the coercive force. The saturation magnetization and coercive force have small values for amorphous bars containing less than 5 at % Al and increase for those containing more than 5 at % Al. The diminishing of the hard magnetic properties for the ribbons and bars having more than 150 μm and 3 mm in thickness, respectively, can be explained by the appearance of the crystalline phases. The hard magnetic behavior of all amorphous samples can be improved by applying thermal treatments at temperatures below the crystallization temperature (between 475 K for alloys without Al and 625 K for Nd_{50}Fe_{40}Al_{10}). The hard magnetic properties of the Nd_{50}Fe_{40}Si_{10−x}Al_{x} alloys are related to the existence of the amorphous phase, which contains ferromagnetic nanoclusters coupled by magnetic exchange interactions.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Nicoleta Lupu

Address for correspondence : National Institute of R&D for technical Physics 47 Mangeron Blvd. 6600 Iasi Romania

Email address : nicole@phys-iasi.ro